OpenStudy (anonymous):

Can Someone Help me Please!?! I need help... -Create and provide the dimensions for two similar figures of your choosing. -What is the similarity ratio of these figures along with the ratio of their surface area and volume? -Show how you go this, either using the actual volumes or using the formula, that the volume ratio is true.

OpenStudy (anonymous):

i would make my life easy and draw two right triangle, say one is 3 - 4 - 5 and the other could be 9 - 12 - 15

OpenStudy (anonymous):

oh that will not work because you want volume, so they have to be three dimensional. so how about two cubes?

OpenStudy (anonymous):

one cube with edge 3 and the other with edge 9 so the ratio of the edges is 3 to 1 , the ration of the volumes is 9^3 to 3^3 or \[\frac{729}{27}=27 \] i.e. 27 to 1

OpenStudy (anonymous):

and the ratio of the surface area is \[\frac{6\times 9^2}{6\times 3^2}=\frac{9^2}{3^2}=9\] i.e. 9 to 1